V. Rates of Return 1. Calculate the dollar rates of return from a £10,000 deposit in a London bank in a year when the interest rate on pounds is 10 percent and the S/E exchange rate moves from $1.50/ to S1.38/E. 2, what would be the real rates of return on the same deposit if there was a simultaneous 10% increase in all dollar prices?
Answer #1

(1)

In current period, Dollar price of deposit = \pounds 10,000 x ($1.50/\pounds) = $15,000

After 1 year, Pound interest = \pounds 10,000 x 10% = \pounds 1,000

After 1 year, Pound value of (Deposit + Interest) = \pounds (10,000 + 1,000) = \pounds 11,000

After 1 year, Dollar value of (Deposit + Interest) = \pounds 11,000 x ($1.38/\pounds) = $15,180

After 1 year, Dollar rate of return = ($15,180 / $15,000) - 1 = 1.012 - 1 = 0.012 = 1.2%

(2)

After 1 year, Nominal Dollar rate of return = 1.2%

After 1 year, Real Dollar rate of return = Nominal Dollar rate of return - Inflation Rate = 1.2% - 10% = -8.8%